# skipabove=0.5skipbelow=0.5align=center

### v1

@@
[…]
skipabove=0.5skipbelow=0.5align=center
Conceptualizing Security Reputation in the Software Supply Chain: A Case for Actor-Based Metrics
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
Conceptualizing Security Reputation in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
A Case for Actor-Based Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
Track Bad Engineers, Not Just Bad Code: Actor-Based Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
A Case for Actor-Based Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
AMSEC: Actor-based metrics for Software Supply Chain Security
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
Repitor: A Case for Actor-Based Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
ARMS: Actor Reputation Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
ARMS: A Vision for Actor Reputation Metrics in the Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
ARMS: A Vision for Actor Reputation Metric Systems in the Open-Source Software Supply Chain
Kelechi G. Kalu1, Sofia Okorafor1, Betül Durak2, Kim Laine2, Radames C. Moreno2,
Santiago Torres-Arias1, and James C. Davis1
1Purdue University
{kalu, sokorafo, santiagotorres, davisjam}@purdue.edu
2Microsoft Research
{betul.durak, kim.laine, radames.cruz}@microsoft.com
Abstract
Many critical information technology and cyber-physical systems rely on a supply chain of open-source software projects. OSS project maintainers often integrate contributions from external actors. While maintainers can assess the correctness of a change request, assessing a change request’s cybersecurity implications is challenging. To help maintainers make this decision, we propose that the open-source ecosystem should incorporate Actor Reputation Metrics (ARMS). This capability would enable OSS maintainers to assess a prospective contributor’s cybersecurity reputation. To support the future instantiation of ARMS, we identify seven generic security signals from industry standards; map concrete metrics from prior work and available security tools, describe study designs to refine and assess the utility of ARMS, and finally weigh its pros and cons.
Index Terms:
Software Supply Chain, Reputation systems
I Introduction
Most commercial software depends on open-source software components [1]. Although this approach reduces development costs, it results in a software supply chain that exposes an organization to cybersecurity risks [2, 3]. Many prior works have examined software supply chain security failures and have developed security techniques [4, 5, 6], engineering processes and frameworks [7, 8, 9]. These works have had substantial success, e.g., the now widely-used Sigstore project for provenance [4], and OpenSSF Scorecard project [10] for process. However, prior works have paid little attention to the actor element of the software supply chain [8].
@@ I Introduction
II Background and Motivation
II-A The Open-Source Software Supply Chain
@@
Open-source software is widely integrated into commercial [11] and government [12] systems. Any individual open-source component is developed by a maintainer team. With their approval, outsiders may be permitted to contribute code [13]. Beyond this direct incorporation of external contributions, each such project often depends on others as components, recursively. This web of interdependencies is a feature of open-source development, allowing (in an idealized world) a reduction in repeated effort [14]. However, each additional point of trust increases the potential attack surface. From the perspective of the downstream application, the result is a software supply chain that can be attacked either through its artifacts or through its actors [15].
@@
We follow the software supply chain definition of Okafor et al. [8]: in their production and distribution, software artifacts undergo a series of operations overseen by actors. This definition indicates that a software supply chain can be secured only through attention to all of these entities.
II-B Artifact-Based Evaluations Are Not Enough
@@
A key activity in open-source projects is expanding the actor pool by introducing new maintainers and contributors into projects [13]. As a software package gains popularity, interest from potential contributors increases [16, 17, 18], often resulting in onboarding new maintainers and merging change requests from new contributors. Evaluating these individuals may involve reputational factors such as community status and connections [19, 20, 16], but security considerations are rarely enforced due to the challenges OSS teams face in managing security resources effectively [21, 22].
@@
As argued by Okafor et al. [8], most cybersecurity work takes an artifact‐based perspective. Efforts to develop static and dynamic security analysis tools (SAST, DAST) [23], refine code reviews [24, 25], and assess artifact provenance [26, 4, 27] are all focused on confirming that an artifact is, to the limits provided by the technique, reliable. While this is not an unreasonable strategy, there are many reasons to avoid relying exclusively on artifact checks, such as:
@@
  • Reliance on known vulnerabilities. Automated scanners typically match code against vulnerability databases; therefore, zero‑day flaws or novel attack vectors often evade detection[28] e.g., Log4j [29].
@@
  • Biased human review. Manual code reviews and audits are applied inconsistently, frequently favoring contributors familiar to project maintainers [19, 18, 30].
@@
  • Scalability constraints. As projects grow, sustaining thorough artifact checks becomes resource‑intensive, leading to superficial reviews or delayed patching [30, 31].
@@
  • Limited socio‑technical insight. Artifact‑centric methods inspect code but ignore the developer behaviors and workflows that often precipitate security issues [32].
@@
II-C Using Actor Reputation to Establish Trust
Reputation systems establish trust between parties who have not been previously connected [33]. When social systems integrate reputation, incentives associated with positive reputation can encourage good behavior over time [34]. Following Hendrikx et al. [35], any reputation system has three interacting entities:
@@ II-C Using Actor Reputation to Establish Trust
  • Trust Engine (Recommender): The broker that supplies the trustor with information about the trustee (interaction data). Its design varies by application and threat model — e.g., for communication [33], online-auction [36], etc.
Two common examples of reputation systems in software engineering are GitHub’s star system [37] and the Stack Overflow point-based system [38]. Both of these systems can be used by a trustor to quantify the number of users satisfied with a trustee’s projects and contributions [39]. As a result, they might influence which GitHub projects an engineer (trustor) may deem to be reliable [40], and which Stack Overflow answers may be trusted by their readers [41].
@@
III Threat Model
III-A Threat Actors
The goal of ARMS is to provide project maintainers with measurements of the security expertise of prospective contributors or maintainers. ARMS operates within a context with three kinds of threat actors:
  1. 1.
    Inexperienced Contributors/Inadvertent Vulnerability: Contributors who lack sufficient security expertise — e.g., they are unfamiliar with standard security practices and tooling within the ecosystem — attempt to join or maintain OSS projects, potentially introducing vulnerabilities through mistakes [42, 43].
  2. 2.
@@ III-A Threat Actors
  3. 3.
    Impersonation. Impersonation occurs when a malicious actor gains control of a legitimate user’s account (e.g., via key compromise [44, 45]).
Our ARMS approach considers the first two classes of threat. The third class, impersonation threats, are out of scope — they undermine the assumption of stable identities necessary for a reputation-based system [46].
III-B Examples
@@
III-B1 Dexcom (Inadvertent Vulnerability)
@@
Dexcom is a medical device company whose products include continuous glucose monitors (CGMs) used by diabetics [47]. Their CGM products were the first to incorporate “smart” capabilities such as pushing health notifications to one’s smartphone. In 2019, Dexcom’s engineers made an error leading to a service outage, resulting in a lack of notifications; many were hospitalized and at least one death is attributable [48]. This was the second such outage in a 12-month window. Although we presume that Dexcom is not intentionally harming its customers, its engineers’ inability to sustain a safety-critical system suggests inadequate experience for this class of work.
III-B2 XZ Utils Backdoor (Reputation Spoofing)
@@
In March 2024, a backdoor was discovered in XZ Utils, an open-source compression tool, which allowed attackers to gain root privileges and execute malicious code on affected systems [49]. The vulnerability was introduced by an actor who had built trust within the project through non-malicious contributions, and they were eventually promoted to co-maintainer [49]. In retrospect, this particular actor had several suspicious reputational signals, most notably having no accounts on other sites and no history of contributions to other projects.
III-B3 ESLint Credential Compromise (Impersonation)
@@
ESLint, a widely used static analysis tool in the npm ecosystem for scanning JavaScript code, was compromised on July 12, 2018 [45]. Attackers gained access to a maintainer’s npm credentials and published malicious package updates to the npm registry [50]. Because the attacker used the identify of a reputable maintainer, a reputation system (which assume stable identities) could not anticipate this attack.
@@
Figure 1: Overview of proposed ARMS system and context case study. Potential contributors (trustees), who may be malicious (Actor A), inadequately expertised (Actor B), or genuine (and capable) (Actor C), express interest or submit change requests . The maintainer team (trustors) requests reputation information on these contributors. The ARMS system retrieves each contributor’s interaction history and quantifies it using the defined security signals and metrics (Interaction data formatter & security signal scoring). Next, the reputation calculator weights these signal values by package usage, community tenure, and centrality, then composites the results and compares them to ecosystem-wide benchmarks (Impact & Benchmark Scoring). Finally, each trustee’s reputation score and recommended action are provided to the maintainer team.
IV ARMS Conceptual Model
To formalize an OSS actor reputation system for GitHub, we propose a reputation system based on the reference model of Hendrikz et al. [35] described earlier, and adapt it to the OSS supply chain context. The following subsections outline our proposed system, define security metrics for an actor’s security reputation, and propose evaluation metrics to measure the effectiveness of these security metrics.
IV-A System Overview
Our proposed system follows the three-element model of Hendrikx et al. (section II-C) comprising a trustor, a trustee, and a trust engine. In the open-source supply chain, the trustor and trustee already exist—e.g., the maintainer team serves as the trustor, and a potential contributor is the trustee. Our work focuses on operationalizing the trust engine component, which is currently absent from the open-source ecosystem.
We describe our proposed system and case study in Figure 1. Interaction history defines the core of our reputation computation, and in our system, we focus specifically on security-related interactions defined by recommended security practices. Although some work has been done on trust establishment in OSS [51], the proposed frameworks are based on defining trust, our work operationalizes trust with reputation systems. In the next section, we define metrics to assess security interactions and history within a typical OSS ecosystem.
Table I: Proposed security signals. Bolded signal categories are derived from security recommendations (e.g, NIST, CNCF, etc). The proposed evaluation metrics in Table II offer ways to evaluate the signals below.
@@ IV-A System Overview
1. Security vulnerability in artifact – Pull Requests/Commits  [51, 52, 53] Measures vulnerabilities introduced through pull requests or commits. Introduced either by the user or to a project owned by the user.
@@
2. Security vulnerability in artifact – use of vulnerable dependencies  [51, 52] Assesses the use and introduction of vulnerable dependencies into the artifact.
@@
3. Use of ecosystem code scanning and security analysis features[54] Evaluates usage of ecosystem’s security tools for code scanning.
4. Use of ecosystem integrity guarantees [27] Examines use of ecosystem integrity features (e.g., code signing).
@@
5. Use of branch protection [55, 6] Checks if branch protection is enforced to prevent unauthorized changes.
@@
6. Use of security policies and vulnerability reporting [55] Determines if the project has security policies and reporting mechanisms.
7. Use of automated workflows [55] Assesses use of [55] automation in workflows to enforce security.
REPUTATION WEIGHTAGE SIGNALS
W1. Package Usage[56] Weighs the utility of the potential contributor’s owned projects
W2. Community Tenure [51] Weighs the the length of the contributor’s belonging to the ecosystem
W3. Centrality Score[18] Weighs the degree of connections of the potential contributor to other actors in the ecosystem.
IV-B Interaction Data – Security Signals and Metrics Definitions

### v2

# ARMS: A Vision for Actor Reputation Metric Systems in the Open-Source Software Supply Chain

@@
Abstract.
Many critical information technology and cyber-physical systems rely on a supply chain of open-source software projects. OSS project maintainers often integrate contributions from external actors. While maintainers can assess the correctness of a pull request, assessing a pull request’s cybersecurity implications is challenging. To help maintainers make this decision, we propose that the open-source ecosystem should incorporate Actor Reputation Metric Systems (ARMS). This capability would enable OSS maintainers to assess a prospective contributor’s cybersecurity reputation. To support the future instantiation of ARMS, we identify seven generic security signals from industry standards; map concrete metrics from prior work and available security tools, describe study designs to refine and assess the utility of ARMS, and finally weigh its pros and cons.
Software Supply Chain, Reputation System
$c$s: Security and privacy Software security engineering
1. Introduction
Most commercial software depends on open-source software components (Sonatype, 2021). Although this approach reduces development costs, it results in a software supply chain that exposes an organization to cybersecurity risks (Benthall, 2017; Willett, 2023). Many prior works have examined software supply chain security failures and have developed security techniques (Newman et al., 2022; Torres-Arias et al., 2019; Zahan et al., 2023), engineering processes and frameworks (do Amaral and Gondim, 2021; Okafor et al., 2022; Melara and Bowman, 2021). These works have had substantial success, e.g., the now widely-used Sigstore project for provenance (Newman et al., 2022), and the OpenSSF Scorecard project (OpenSSF, [n. d.]) for process. However, prior works have paid little attention to the actor element of the software supply chain (Okafor et al., 2022), which we believe is now the weaker link.
@@ 1. Introduction
  • We motivate actor-centric reputation as a complementary lens to artifact-centric security measurement, using recent supply-chain incidents and maintainer workflows to highlight gaps in current vetting practices.
@@
Workshop positioning & Next Steps. This submission describes our vision and system design. We seek feedback from the JAWS community on (i) the proposed experiments for assessing the suitability and effectiveness of our actor-centric security signals, (ii) appropriate evaluation targets and baselines, and (iii) practical deployment considerations, including fairness for newcomers and robustness to missing or private data.
2. Background & Motivation
2.1. The Open-Source Software Supply Chain
@@
Open-source software is widely integrated into commercial (Franke et al., 2024) and government (U.S. Department of Defense, 2009) systems. Any individual open-source component is developed by a maintainer team. With their approval, outsiders may be permitted to contribute code (Raymond, 1999). Beyond this direct incorporation of external contributions, each such project often depends on others as components, recursively. This web of interdependencies is a feature of open-source development, allowing (in an idealized world) a reduction in repeated effort (Sweeney, 2023). However, each additional point of trust increases the potential attack surface. From the perspective of the downstream application, the result is a software supply chain that can be attacked either through its artifacts or through its actors (Ohm et al., 2020).
@@
We follow the software supply chain definition of Okafor et al. (Okafor et al., 2022): in their production and distribution, software artifacts undergo a series of operations overseen by actors. This definition indicates that a software supply chain can be secured only through attention to all of these entities.
2.2. Artifact-Based Evaluations Are Not Enough
@@
A key activity in open-source projects is expanding the actor pool by introducing new maintainers and contributors into projects (Raymond, 1999). As a software package gains popularity, interest from potential contributors increases (Maldeniya et al., 2020; Qiu et al., 2019; Hamer et al., 2025), often resulting in onboarding new maintainers and merging pull requests from new contributors. Evaluating these individuals may involve reputational factors such as community status and connections (Tsay et al., 2014; Yu et al., 2014; Maldeniya et al., 2020), but security considerations are rarely enforced due to the challenges OSS teams face in managing security resources effectively (Wermke et al., 2022; Amft et al., 2024).
@@
As argued by Okafor et al. (Okafor et al., 2022), most cybersecurity work takes an artifact‐based perspective. Efforts to develop static and dynamic security analysis tools (SAST, DAST) (Elder et al., 2022), refine code reviews (Howard, 2006; Braz and Bacchelli, [n. d.]), and assess artifact provenance (Vu et al., 2021; Newman et al., 2022; Schorlemmer et al., 2024) are all focused on confirming that an artifact is, to the limits provided by the technique, reliable. While this is not an unreasonable strategy, there are many reasons to avoid relying exclusively on artifact checks, such as:
@@
  • Reliance on known vulnerabilities. Automated scanners typically match code against vulnerability databases; therefore, zero‑day flaws or novel attack vectors often evade detection(Pan et al., 2024) e.g., Log4j (Hiesgen et al., [n. d.]).
@@
  • Biased human review. Manual code reviews and audits are applied inconsistently, frequently favoring contributors familiar to project maintainers (Tsay et al., 2014; Hamer et al., 2025; Kononenko et al., 2016).
@@
  • Scalability constraints. As projects grow, sustaining thorough artifact checks becomes resource‑intensive, leading to superficial reviews or delayed patching (Kononenko et al., 2016; Harman and O’Hearn, 2018).
@@
  • Limited socio‑technical insight. Artifact‑centric methods inspect code but ignore the developer behaviors and workflows that often precipitate security issues (Kalu et al., 2023).
@@
2.3. Using Actor Reputation to Establish Trust
Reputation systems establish trust between parties who have not been previously connected (Durak et al., 2023). When social systems integrate reputation, incentives associated with positive reputation can encourage good behavior over time (Josang et al., 2007). Following Hendrikx et al. (Hendrikx et al., 2015), any reputation system has three interacting entities:
@@ 2.3. Using Actor Reputation to Establish Trust
  • Trust Engine (Recommender): The broker that supplies the trustor with information about the trustee (interaction data). Its design varies by application and threat model — e.g., for communication (Durak et al., 2023), online-auction (Goodrich and Kerschbaum, 2011), etc.
Two common examples of reputation systems in software engineering are GitHub’s star system (GitHub Docs, [n. d.]) and the Stack Overflow point-based system (Stack Overflow Help Center, [n. d.]). Both of these systems can be used by a trustor to quantify the number of users satisfied with a trustee’s projects and contributions (Cameron, 2005). As a result, they might influence which GitHub projects an engineer (trustor) may deem to be reliable (Borges and Valente, 2018), and which Stack Overflow answers may be trusted by their readers (Wang et al., 2021).
@@
3. Threat Model
3.1. Threat Actors
The goal of ARMS is to provide project maintainers with measurements of the security expertise of prospective contributors or maintainers. ARMS operates considers three kinds of threat actors:
  1. (1)
    Inexperienced Contributors/Inadvertent Vulnerability: Contributors who lack sufficient security expertise — e.g., they are unfamiliar with standard security practices and tooling within the ecosystem — attempt to join or maintain OSS projects, potentially introducing vulnerabilities through mistakes (Norman, 2013; Reason, 1990).
  2. (2)
@@ 3.1. Threat Actors
  3. (3)
    Impersonation. Impersonation occurs when a malicious actor gains control of a legitimate user’s account (e.g., via key compromise (Codecov Security Team, 2021; ESLint Team, 2018)).
Our ARMS approach considers the first two classes of threat. The third class, impersonation threats, are out of scope — they undermine the assumption of stable identities necessary for a reputation-based system (Douceur, 2002).
3.2. Examples
@@
3.2.1. Dexcom (Inadvertent Vulnerability)
@@
Dexcom is a medical device company whose products include continuous glucose monitors (CGMs) used by diabetics (Mukherjee, 2019). Their CGM products were the first to incorporate “smart” capabilities such as pushing health notifications to one’s smartphone. In 2019, Dexcom’s engineers made an error leading to a service outage, resulting in a lack of notifications; many were hospitalized and at least one death is attributable (O’Connor, 2019). This was the second such outage in a 12-month window. Although we presume that Dexcom is not intentionally harming its customers, its engineers’ inability to sustain a safety-critical system suggests inadequate experience for this class of work.
3.2.2. XZ Utils Backdoor (Reputation Spoofing)
@@
In March 2024, a backdoor was discovered in XZ Utils, an open-source compression tool, which allowed attackers to gain root privileges and execute malicious code on affected systems (Lins et al., 2024). The vulnerability was introduced by an actor who built trust within the project through non-malicious contributions, and they were eventually promoted to co-maintainer (Lins et al., 2024). In retrospect, this actor showed suspicious reputational signals, including limited public contribution history outside the target project and a narrow set of early contributions.
3.2.3. ESLint Credential Compromise (Impersonation)
@@
ESLint, a widely used static analysis tool in the npm ecosystem for scanning JavaScript code, was compromised on July 12, 2018 (ESLint Team, 2018). Attackers gained access to a maintainer’s npm credentials and published malicious package updates to the npm registry (Cycode Security Team, 2022). Because the attacker used the identify of a reputable maintainer, a reputation system (which assume stable identities) could not anticipate this attack.
@@
Figure 1. Overview of the proposed ARMS system and context case study. Potential contributors (trustees), who may be malicious (Actor A), inexperienced (Actor B), or genuine (and capable) (Actor C), express interest or submit pull requests . The maintainer team (trustors) requests reputation information on these contributors. The ARMS system retrieves each contributor’s interaction history and quantifies it using the defined security signals and metrics (Interaction data formatter & security signal scoring). Next, the reputation calculator calibrates these signal values by package usage, community tenure, and centrality, then composites the results and compares them to ecosystem-wide benchmarks (Impact & Benchmark Scoring). Finally, each trustee’s reputation score and recommended action are provided to the maintainers.
Table 1. Proposed security signals, weighting factors, and evaluation metrics. Bolded security and weighting signals are derived from established frameworks (e.g., NIST, SLSA, CNCF) and prior work, respectively. The proposed metrics describe how to measure each highlighted security or weighting signal. Relevant case studies and related research are cited to justify inclusion of select metrics. Note: S3–S7 reflect repository configuration and are treated as context unless governance control is attributable to the actor (e.g., owner/maintainer).
@@
1. Security vulnerability in artifact: Pull Requests/Commits Measures vulnerabilities introduced through pull requests or commits, either by the user or in projects owned by the user.
a. Time to fix/close security vulnerabilities of various severity levels (Boughton et al., 2024) Time taken to address vulnerabilities detected in pull requests or commits.
b. Severity levels of security vulnerability reported (Shukla et al., 2019) Evaluates the criticality of reported vulnerabilities (low, medium, high) for the user’s projects.
c. Presence of vulnerability in PR/commits (not linked to external dependencies) (Gonzalez et al., 2021) Measures vulnerabilities that stem from the user’s direct contributions.
@@
2. Security vulnerability in artifact: use of vulnerable dependencies Assesses the use and introduction of vulnerable dependencies into the artifact.
a. Length of time before fix (Boughton et al., 2024) Time taken to resolve vulnerable dependencies after detection (or public disclosure).
b. Severity levels of Repos/Project’s reported vulnerability (Shukla et al., 2019) Tracks severity of vulnerabilities in dependencies.
c. Number of Repos/Projects that have a vulnerable dependency Total number of projects affected by vulnerable dependencies as a percentage of the user’s total projects.
d. Number of vulnerable dependencies per project Assesses how widespread vulnerable dependencies are within each project.
e. Number of projects with reported vulnerable dependencies Measures the overall exposure of the user’s projects to vulnerable dependencies.
Repository governance security signals (S3–S7)
@@
3. Use of ecosystem code scanning and security analysis features Evaluates usage of the ecosystem’s security tools for code scanning.
a. Status of dependabot alerts (for dependencies) (Fischer et al., 2023) Monitors whether security alerts for dependencies are being addressed.
b. Status of Secret sharing/Validity Checks Evaluates management of secret scanning and validity checks.
c. Code scanning and Vulnerability alerts (Fischer et al., 2023) Assesses whether security scanning tools are actively used and vulnerability alerts managed.
d. Push Protection Status Checks whether protection features are used to block pushes with exposed secrets.
a. Number of Projects with an integrity guarantee (Schorlemmer et al., 2024) Measures how many projects are secured with integrity verification features such as code signing.
@@
5. Use of branch protection Checks whether branch protection is enforced to prevent unauthorized changes.
a. Number of protected branches per project Evaluates how many branches within each project are guarded against direct commits.
b. Number of Projects with Protected branches (OpenSSF, [n. d.]; Zahan et al., 2023) Tracks how many projects enforce branch protection.
@@
6. Use of security policies and vulnerability reporting Determines whether the project has security policies and reporting mechanisms.
a. Status of private vulnerability reporting, or security policy (OpenSSF, [n. d.]) Checks whether private vulnerability reporting channels and security policies are established and functional.
7. Use of automated workflows Assesses use of automation in workflows to enforce security.
a. Number of Projects with Automated workflows (OpenSSF, [n. d.]) Measures how many user projects use automated workflows for security enforcement (e.g., required checks, automated tests).
CALIBRATION FACTORS
W1. Package Usage (Boysel, 2023)
a. Number of downloads/stars/forks Calibrates (Weights) project-level evidence by downstream exposure.
W2. Community Tenure (Boughton et al., 2024)
a. Length of time contributing to project Calibrates uncertainty based on time contributing to other projects.
b. Length of time owning account Calibrates uncertainty based on account age.
c. Contributory Strength Calibrates evidence strength based on contribution volume (e.g., commits, issues, pull requests).
W3. Centrality Score (Hamer et al., 2025)
a. Connections to other Actors (Hamer et al., 2025) Calibrates ecosystem context based on a contributor’s network of interactions.
b. Time to form connections Calibrates based on how long it took to establish the connections in (a).
4. ARMS Conceptual Model
To formalize an OSS actor reputation system for GitHub, we propose a reputation system based on the reference model of Hendrikz et al. (Hendrikx et al., 2015) described earlier, and adapt it to the OSS supply chain context. The following subsections outline our proposed system, define signals for an actor’s security reputation, and propose computations to operationalize these signals in practice.
4.1. System Overview
Our proposed system follows the three-element model of Hendrikx et al. (section 2.3) comprising a trustor, a trustee, and a trust engine. In the open-source supply chain, the trustor and trustee already exist—e.g., the maintainer team serves as the trustor, and a potential contributor is the trustee. Our work focuses on operationalizing the trust engine component, which is currently absent from the open-source ecosystem.
We describe our proposed system in Figure 1. Interaction history defines the core of our reputation computation, and in our system, we focus specifically on security-related interactions defined by recommended security practices. Although some work has been done on trust establishment in OSS (Boughton et al., 2024), the proposed frameworks are based on defining trust, our work operationalizes trust with reputation systems. In the next section, we define metrics to assess security interactions and history within a typical OSS ecosystem.
4.2. Interaction Data – Security Signals and Metrics Definitions